Could you help me with the following problem?
Solve the system algebraically.
5x+2y=10
3x+2y=6
Make sure there are no spaces in your answer. Be sure to include a comma.
Answer{( )}

Answer by RAY100(1637)

1) 5x +2y = 10
2) 3x +2y =6
subt (1) - (2)
2x =4
divide by 2
x=2
subst in (1)
5(2) +2y =10
10 +2y =10
subt 10 both sides
2y =0
y=0
answer (2,0)
checking
1) 5(2) +2(0) =10 ok
2) 3(2) = 2(0) =6 ok